# Why Does Little Robustness Help? A Further Step Towards Understanding Adversarial Transferability

### v7

@@ 1 Introduction
Adversarial transferability is an intriguing property of adversarial examples (AEs), where examples crafted against a surrogate DNN could fool other DNNs as well. Various techniques [45, 13, 24, 12, 32, 86, 76, 88, 26] have been proposed for the generation process of AEs to increase transferability [^1] We refer to “adversarial transferability” as “transferability”., such as integrating momentum to stabilize the update direction [12] or applying transformations at each iteration to create diverse input patterns [76]. At a high level, all these research endeavors to find transferable AEs under given surrogate models by proposing complex tricks to be incorporated into the generation pipeline one after another. However, this approach could result in non-trivial computational expenses and low scalability [49].
@@
With a different perspective, another line of works [22, 23, 72, 34] start to examine the role of surrogate models. One useful observation is that attacking an ensemble of surrogate models with different architectures can obtain a more general update direction for AEs [38, 5]. A recent study [23] also proposed fine-tuning a well-trained surrogate to obtain a set of intermediate models that can be used for an ensemble. However, it remains unclear which type of surrogate performs the best and should be included in the ensemble. Our work taps into this line and tries to manipulate surrogates to launch stronger transfer attacks.
@@
Meanwhile, Demontis et al. [9] started to explain the transferability property and showed that model complexity and gradient alignment negatively and positively correlate with transferability, respectively. Lately, Yang et al. [78] established a transferability lower bound and theoretically connect transferability to two key factors: model smoothness and gradient similarity. Model smoothness captures the general invariance of the gradient w.r.t. input features for a given model, while gradient similarity refers to the alignment of the gradient direction between surrogate and target models. However, there is still much to be understood and explored regarding these factors. For instance, it is unclear which factor plays a more important role in regulating transferability, how existing empirical findings for improving transferability affect these factors, and how to generally optimize them both simultaneously. Such a complex situation makes it challenging to completely understand the role of surrogates. Consequently, there is currently no consensus on how to construct better surrogates to achieve higher adversarial transferability. In support of this, a recent work [42], after performing a large-scale empirical study in real-world settings, concluded that surrogate-level factors that affect the transferability highly chaotically interact and there is no effective solution to obtain a good surrogate except through trial and error.
@@
Stronger regularized (smoother) models provide better surrogates on average [9]. (1) AT with large budget yields smoother models that degrade transferability. In Sec. 2.1. (2) Stronger regularizations cannot always outperform less smooth solutions like SAM. In Sec. 4.3. Partly  conflicting
AT and data augmentation do not show strong correlations to transfer attacks in the “real-world” environment [42]. (1) AT with small budget benefits transfer attack while large budget hinders it. (2) Data augmentation generally impairs transfer attacks, especially for stronger augmentations. In Sec. 6, Q6. Conflicting
Surrogate models with better generalization performance could result in more transferable AEs [68]. Data augmentations that yield surrogates with the best generalization perform the worst in transfer attacks. In Sec. 6, Q4. Conflicting
Attacking multiple surrogates from a sufficiently large geometry vicinity (LGV) benefits transferability [23]. Attacking multiple surrogates from arbitrary LGV of a single superior surrogate may degrade transferability. In Sec. 5.2. Partly  conflicting
Regularizing pressure transfers from the weight space to the input space. [11]. This transfer effect exists, yet is marginal and unstable. In Sec. 4.2. Partly  conflicting
The poor transferability of ViT is because existing attacks are not strong enough to fully exploit its potential [46]. The transferability of ViT may have been restrained by its default training paradigm. In Sec. 6, Q5 Parallel
Model complexity (the number of local optima in loss surface) correlates with transferability [9]. A smoother model is expected to have less and wider local optima in a finite space. In Sec. 6, Q3 Causal
AEs lie off the underlying manifold of clean data [21]. Adversarial training causes data distribution shift induced by off-manifold AEs, thus impairing gradient similarity. Dependent
(1) Attacking an ensemble of surrogates in the distribution found by Bayes learning improves the transferability [33]. (2) SAM can be seen as a relaxation of Bayes [44]. SAM yields general input gradient alignment towards every training solution. Attacking SAM solution significantly improves transferability. In Sec. 6, Q7. Matching
@@
Gaining a deeper understanding of the “little robustness” phenomenon. We begin by exploring the intriguing “little robustness” phenomenon [57], where models produced by small-budget adversarial training (AT) serve as better surrogates than standard ones (see Fig. 2). Through tailored definitions of gradient similarity and model smoothness, we recognize this phenomenon comes along with an inherent trade-off effect between these two factors on transferability. Thus we attribute the “little robustness” appearance to the persistent improvement of model smoothness and deterioration of gradient similarity. Further, we identify the observed gradient dissimilarity in AT as the result of the data distribution shift caused by off-manifold samples.
@@ 2 Explaining Little Robustness
Understanding adversarial transferability is still a challenging task, and many empirical observations in the literature are rather perplexing. As a typical case, “little robustness” indicates that a model adversarially trained with small perturbation budgets improves the transferability [57, 58]. It remains unclear why small-budget adversarially trained models can serve as better surrogates, whereas large-budget ones do not exhibit this benefit. In this section, we tailor the definitions of gradient similarity and model smoothness, two recently proposed concepts that are believed to be important for transferability [78], to analyze this phenomenon.
@@ 2.1 Transferability Circuit of Adversarial Training
Adversarial training [40] (AT) uses adversarial examples generated on $\mathbf{S}$ as augmented data and minimizes the following adversarial loss:
@@
where $\delta\in\mathbb{R}^{d}$ denotes the adversarial perturbation and $\epsilon$ is the adversarial budget. As evidenced by our experiments in Fig. 2, the attack success rates (ASRs) rise with relatively small-budget adversarial trained models and start to decrease with the high-budget ones. This “transferability circuit” is particularly intriguing. We thus revisit the recently proposed transferability lower bound [78] and utilize it as a tool to analyze the underlying reason.
@@ 2.2 Lower Bound of Transferability
A smoother model on $\mathcal{D}$ is featured with smaller $\sigma_{\mathcal{F}}$ and $\overline{\sigma}_{\mathcal{F}}$, indicating more invariance of loss gradient. Different from its original definition where the global Lipschitz constant $\beta_{\mathcal{F}}$ is used [78], we use this curvature metric (i.e., the dominant eigenvalue of the Hessian) to define the model smoothness in a local manner. With this modified definition, we can theoretically and empirically quantify the smoothness of loss surface on $\mathcal{F}$ for a given data distribution $\mathcal{D}$. Moreover, this provides a tighter bound for the model’s loss function gradient, namely, an explicit relation $\overline{\sigma}_{\mathcal{F}}{\color[rgb]{0,0,0}\leq}\beta_{\mathcal{F}}$ holds. Although the true distribution $\mathcal{D}$ is unknown, one can still sample a set of data from $\mathbf{S}$ and use the empirical mean as an approximation to evaluate the model smoothness. Note that our definition is correlated to the model complexity concept introduced in Demontis et al. [9] (see Sec. 6, Q3).
@@
Given a distribution $\mathcal{D}$, we can further define the infimum of loss gradient similarity on $\mathcal{D}$ as $\underline{\mathcal{S}}{\color[rgb]{0,0,0}{}_{\mathcal{D}}}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}}\right)=\inf\limits_{(x,y)\sim\mathcal{D}}\mathcal{S}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}},x,y\right)$. Similar to Demontis et al. [9], we define the expected loss gradient similarity as $\tilde{\mathcal{S}}{\color[rgb]{0,0,0}{}_{\mathcal{D}}}(\ell_{\mathcal{F}},\ell_{\mathcal{G}})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{S}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}},x,y\right)]$, which can capture the similarity between two models on $\mathcal{D}$. Naturally, a larger gradient similarity indicates a more general alignment between the adversarial directions of two given models. We can sample a set of data from $\mathbf{S}$ to evaluate this alignment between two models on $\mathcal{D}$.
@@
Here we define transferability the same way as [78] for untargeted attacks at the instance level. A successful untargeted transfer attack requires both the surrogate and the target models to give correct predictions for the unperturbed input and incorrect predictions for the perturbed one. The targeted version is similar. Note that we abuse the notations $\delta$ and $\epsilon$ in Definitions 3 and 1, while the notations in Eq. 1 have similar meanings under different contexts.
@@
$c_{\mathcal{G}}$ $=\max_{(x,y)\sim\mathcal{D}}\frac{\min\limits_{y^{\prime}\in\mathcal{L}:y^{\prime}\neq y}\ell_{\mathcal{G}}\left(x+\delta,y^{\prime}\right)-\ell_{\mathcal{G}}(x,y)+\overline{\sigma}_{\mathcal{G}}\epsilon^{2}/2}{\left\|\nabla_{x}\ell_{\mathcal{G}}(x,y)\right\|_{2}}.$
The natural risks (inaccuracy) of models $\mathcal{F}$ and $\mathcal{G}$ on $\mathcal{D}$ are defined as $\gamma_{\mathcal{F}}$ and $\gamma_{\mathcal{G}}$ respectively, and $\overline{\sigma}_{\mathcal{F}}$ and $\overline{\sigma}_{\mathcal{G}}$ are the upper smoothness of $\mathcal{F}$ and $\mathcal{G}$.
Note that we deliver the lower bound of transferability for untargeted attacks as an example, and the targeted attack follows a similar form. The conclusions for the targeted attack agree with the untargeted case, as suggested in [78], so we omit it for space. The proof is provided in the Appendix. Although the theorem is initially for $l_{2}$-norm perturbations, it can be generalized to $l_{\infty}$-norm as demonstrated in [78].
Implications. Different from [78], which seeks to suppress the transferability from both surrogate and target sides, this paper focuses on the surrogate aspect alone to increase transferability. Therefore, we next briefly analyze how all the terms relating to $\mathcal{F}$ influence the lower bound.
@@
  • natural risk $\gamma_{\mathcal{F}}$: the negative effect of natural risk on the lower bound of transferability is obvious. A model with lame accuracy on $\mathcal{D}$ is certainly undesirable for adversarial attacks. Consequently, we minimize this term by training models with sufficient epochs for CIFAR-10 and fine-tuning pre-trained ImageNet classifiers for ImageNette.
@@
It is worth noting that a sensible adversary naturally desires to use more precise surrogates and stronger attacks against unknown targets. However, smoothness and similarity are more complex to understand and optimize compared to fooling probability and natural risk. Therefore, we focus on investigating the connection between smoothness, similarity, and transferability as well as how the various training mechanisms regulate them under the restriction that the other two factors above are fairly acceptable. Prior study [9] has shown their link with transferability separately, while we explore their joint effect on transferability.
@@ 2.3 Trade-off Under Adversarial Training
Table II: Untargeted and targeted transfer ASRs of AEs crafted against baseline (ST), data augmentations, adversarial training, and gradient regularizations surrogates under the different $L_{\infty}$ budgets (4/255, 8/255, 16/255) on CIFAR-10 using PGD. Results in gray are (more than 0.2%) below the baseline, results in white are close to the baseline (±$\pm 0.2\%$). Results in darkest orange are the highest in each respective column, the second darkest are the second highest, and the least dark are others more than 0.2% above the baseline. We conduct experiments using surrogate models trained from 3 different random seeds in each setting and report the average results and the according error bars. In %.
@@
Data distribution shift impairs gradient similarity. Different from model smoothness, it is difficult to understand how exactly AT degrades the gradient similarity. In this work, we attribute this degradation to the data distribution shift induced by the off-manifold examples in AT. Since the emerging of adversarial examples [61], there is a long-held belief that: Clean data lies in a low-dimensional manifold. Even though the adversarial examples are close to the clean data, they lie off the underlying data manifold [39, 56, 30, 21, 37]. Nevertheless, recent researches demonstrate that AEs can also be on-manifold [31, 48, 59, 36, 37], and on-manifold and off-manifold AEs may co-exist [75]. With this in mind, we analyze how AT enlarges the data distribution shift along with the increment of adversarial budget $\epsilon$.
Given a regular loss function $\ell(f(x),y)$ on the low-dimensional manifold $\mathcal{M}$, the adversarial perturbation $\delta\in\mathbb{R}^{d}$ could make the loss $\ell(f(x+\delta),y)$ sufficiently high with $x+\delta\in\mathcal{M}$ or $x+\delta\notin\mathcal{M}$ [77]. At a high level, AT augments the dataset by adding adversarial examples during each iteration, obtaining an augmented data distribution on $\mathcal{P}_{\mathcal{X}}$, and the adversarial budget $\epsilon$ can be regarded as a parameter that controls the augmentation magnitude. Intuitively, a larger $\epsilon$ induces more space for off-manifold samples, resulting in a larger distribution shift. On the other hand, it is also widely acknowledged that adversarial training with larger perturbation budgets leads to robust overfitting [73, 14, 60, 50]. Accordingly, a larger $\epsilon$ may cause more disbenefit of underfitting to the normal distribution $\mathcal{D}$. Formally, we formalize our hypothesis as follows:

### v8

@@ 1 Introduction
Adversarial transferability is an intriguing property of adversarial examples (AEs), where examples crafted against a surrogate DNN could fool other DNNs as well. Various techniques [Naseer_2021_ICCV, Dong_2019_CVPR, NEURIPS2020_00e26af6, Dong_2018_CVPR, Li_2020_CVPR, zhao2021on, xie2019improving, zhou2023advencoder, hu2022protecting] have been proposed for the generation process of AEs to increase transferability [^1] We refer to “adversarial transferability” as “transferability”., such as integrating momentum to stabilize the update direction [Dong_2018_CVPR] or applying transformations at each iteration to create diverse input patterns [xie2019improving]. At a high level, all these research endeavors to find transferable AEs under given surrogate models by proposing complex tricks to be incorporated into the generation pipeline one after another. However, this approach could result in non-trivial computational expenses and low scalability [qin2022boosting].
@@
With a different perspective, another line of works [pmlr-v180-gubri22a, gubriLGV2022, Wu2020Skip, li2020learning] start to examine the role of surrogate models. One useful observation is that attacking an ensemble of surrogate models with different architectures can obtain a more general update direction for AEs [liu2017delving, chen2023rethinking]. A recent study [gubriLGV2022] also proposed fine-tuning a well-trained surrogate to obtain a set of intermediate models that can be used for an ensemble. However, it remains unclear which type of surrogate performs the best and should be included in the ensemble. Our work taps into this line and tries to manipulate surrogates to launch stronger transfer attacks.
@@
Meanwhile, Demontis et al. [security19] started to explain the transferability property and showed that model complexity and gradient alignment negatively and positively correlate with transferability, respectively. Lately, Yang et al. [yangNeurIPS2021TRS2021] established a transferability lower bound and theoretically connect transferability to two key factors: model smoothness and gradient similarity. Model smoothness captures the general invariance of the gradient w.r.t. input features for a given model, while gradient similarity refers to the alignment of the gradient direction between surrogate and target models. However, there is still much to be understood and explored regarding these factors. For instance, it is unclear which factor plays a more important role in regulating transferability, how existing empirical findings for improving transferability affect these factors, and how to generally optimize them both simultaneously. Such a complex situation makes it challenging to completely understand the role of surrogates. Consequently, there is currently no consensus on how to construct better surrogates to achieve higher adversarial transferability. In support of this, a recent work [maoP2022TransferAttacks2022], after performing a large-scale empirical study in real-world settings, concluded that surrogate-level factors that affect the transferability highly chaotically interact and there is no effective solution to obtain a good surrogate except through trial and error.
@@
Stronger regularized (smoother) models provide better surrogates on average [security19]. (1) AT with large budget yields smoother models that degrade transferability. In Sec. 2.1. (2) Stronger regularizations cannot always outperform less smooth solutions like SAM. In Sec. 4.3. Partly conflicting
AT and data augmentation do not show strong correlations to transfer attacks in the “real-world” environment [maoP2022TransferAttacks2022]. (1) AT with small budget benefits transfer attack while large budget hinders it. (2) Data augmentation generally impairs transfer attacks, especially for stronger augmentations. In Sec. 6, Q6. Conflicting
Surrogate models with better generalization performance could result in more transferable AEs [wang2022role]. Data augmentations that yield surrogates with the best generalization perform the worst in transfer attacks. In Sec. 6, Q4. Conflicting
Attacking multiple surrogates from a sufficiently large geometry vicinity (LGV) benefits transferability [gubriLGV2022]. Attacking multiple surrogates from arbitrary LGV of a single superior surrogate may degrade transferability. In Sec. 5.2. Partly conflicting
Regularizing pressure transfers from the weight space to the input space. [Dherin2022why]. This transfer effect exists, yet is marginal and unstable. In Sec. 4.2. Partly conflicting
The poor transferability of ViT is because existing attacks are not strong enough to fully exploit its potential [naseer2022on]. The transferability of ViT may have been restrained by its default training paradigm. In Sec. 6, Q5 Parallel
Model complexity (the number of local optima in loss surface) correlates with transferability [security19]. A smoother model is expected to have less and wider local optima in a finite space. In Sec. 6, Q3 Causal
AEs lie off the underlying manifold of clean data [gilmer2018adversarial]. Adversarial training causes data distribution shift induced by off-manifold AEs, thus impairing gradient similarity. Dependent
(1) Attacking an ensemble of surrogates in the distribution found by Bayes learning improves the transferability [li2023making]. (2) SAM can be seen as a relaxation of Bayes [mollenhoff2023sam]. SAM yields general input gradient alignment towards every training solution. Attacking SAM solution significantly improves transferability. In Sec. 6, Q7. Matching
@@
Gaining a deeper understanding of the “little robustness” phenomenon. We begin by exploring the intriguing “little robustness” phenomenon [springer2021little], where models produced by small-budget adversarial training (AT) serve as better surrogates than standard ones (see Fig. 2). Through tailored definitions of gradient similarity and model smoothness, we recognize this phenomenon comes along with an inherent trade-off effect between these two factors on transferability. Thus we attribute the “little robustness” appearance to the persistent improvement of model smoothness and deterioration of gradient similarity. Further, we identify the observed gradient dissimilarity in AT as the result of the data distribution shift caused by off-manifold samples.
@@ 2 Explaining Little Robustness
Understanding adversarial transferability is still a challenging task, and many empirical observations in the literature are rather perplexing. As a typical case, “little robustness” indicates that a model adversarially trained with small perturbation budgets improves the transferability [springer2021little, springer2021uncovering]. It remains unclear why small-budget adversarially trained models can serve as better surrogates, whereas large-budget ones do not exhibit this benefit. In this section, we tailor the definitions of gradient similarity and model smoothness, two recently proposed concepts that are believed to be important for transferability [yangNeurIPS2021TRS2021], to analyze this phenomenon.
@@ 2.1 Transferability Circuit of Adversarial Training
Adversarial training [madry2018towards] (AT) uses adversarial examples generated on $\mathbf{S}$ as augmented data and minimizes the following adversarial loss:
@@
where $\delta\in\mathbb{R}^{d}$ denotes the adversarial perturbation and $\epsilon$ is the adversarial budget. As evidenced by our experiments in Fig. 2, the attack success rates (ASRs) rise with relatively small-budget adversarial trained models and start to decrease with the high-budget ones. This “transferability circuit” is particularly intriguing. We thus revisit the recently proposed transferability lower bound [yangNeurIPS2021TRS2021] and utilize it as a tool to analyze the underlying reason.
@@ 2.2 Lower Bound of Transferability
A smoother model on $\mathcal{D}$ is featured with smaller $\sigma_{\mathcal{F}}$ and $\overline{\sigma}_{\mathcal{F}}$, indicating more invariance of loss gradient. Different from its original definition where the global Lipschitz constant $\beta_{\mathcal{F}}$ is used [yangNeurIPS2021TRS2021], we use this curvature metric (i.e., the dominant eigenvalue of the Hessian) to define the model smoothness in a local manner. With this modified definition, we can theoretically and empirically quantify the smoothness of loss surface on $\mathcal{F}$ for a given data distribution $\mathcal{D}$. Moreover, this provides a tighter bound for the model’s loss function gradient, namely, an explicit relation $\overline{\sigma}_{\mathcal{F}}{\color[rgb]{0,0,0}\leq}\beta_{\mathcal{F}}$ holds. Although the true distribution $\mathcal{D}$ is unknown, one can still sample a set of data from $\mathbf{S}$ and use the empirical mean as an approximation to evaluate the model smoothness. Note that our definition is correlated to the model complexity concept introduced in Demontis et al. [security19] (see Sec. 6, Q3).
@@
Given a distribution $\mathcal{D}$, we can further define the infimum of loss gradient similarity on $\mathcal{D}$ as $\underline{\mathcal{S}}{\color[rgb]{0,0,0}{}_{\mathcal{D}}}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}}\right)=\inf\limits_{(x,y)\sim\mathcal{D}}\mathcal{S}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}},x,y\right)$. Similar to Demontis et al. [security19], we define the expected loss gradient similarity as $\tilde{\mathcal{S}}{\color[rgb]{0,0,0}{}_{\mathcal{D}}}(\ell_{\mathcal{F}},\ell_{\mathcal{G}})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{S}\left(\ell_{\mathcal{F}},\ell_{\mathcal{G}},x,y\right)]$, which can capture the similarity between two models on $\mathcal{D}$. Naturally, a larger gradient similarity indicates a more general alignment between the adversarial directions of two given models. We can sample a set of data from $\mathbf{S}$ to evaluate this alignment between two models on $\mathcal{D}$.
@@
Here we define transferability the same way as [yangNeurIPS2021TRS2021] for untargeted attacks at the instance level. A successful untargeted transfer attack requires both the surrogate and the target models to give correct predictions for the unperturbed input and incorrect predictions for the perturbed one. The targeted version is similar. Note that we abuse the notations $\delta$ and $\epsilon$ in Definitions 3 and 1, while the notations in Eq. 1 have similar meanings under different contexts.
@@
$c_{\mathcal{G}}$ $=\max_{(x,y)\sim\mathcal{D}}\frac{\min\limits_{y^{\prime}\in\mathcal{L}:y^{\prime}\neq y}\ell_{\mathcal{G}}\left(x+\delta,y^{\prime}\right)-\ell_{\mathcal{G}}(x,y)+\overline{\sigma}_{\mathcal{G}}\epsilon^{2}/2}{\left\|\nabla_{x}\ell_{\mathcal{G}}(x,y)\right\|_{2}},$
and $\epsilon$ be sufficiently large such that $\epsilon>c_{\mathcal{G}}$ and $\epsilon>c_{\mathcal{F}}$. $\gamma_{\mathcal{F}}$ and $\gamma_{\mathcal{G}}$ are the natural inaccuracies of $\mathcal{F}$ and $\mathcal{G}$ on $\mathcal{D}$, $\overline{\sigma}_{\mathcal{F}}$ and $\overline{\sigma}_{\mathcal{G}}$ are the upper-bound of smoothness w.r.t. $\mathcal{F}$ and $\mathcal{G}$.
Note that we deliver the lower bound of transferability for untargeted attacks as an example, and the targeted attack follows a similar form. The conclusions for the targeted attack agree with the untargeted case, as suggested in [yangNeurIPS2021TRS2021], so we omit it for space. The proof is provided in the Appendix. Although the theorem is initially for $l_{2}$-norm perturbations, it can be generalized to $l_{\infty}$-norm as demonstrated in [yangNeurIPS2021TRS2021].
Implications. Different from [yangNeurIPS2021TRS2021], which seeks to suppress the transferability from both surrogate and target sides, this paper focuses on the surrogate aspect alone to increase transferability. Therefore, we next briefly analyze how all the terms relating to $\mathcal{F}$ influence the lower bound.
@@
  • natural inaccuracy $\gamma_{\mathcal{F}}$: the negative effect of natural inaccuracy on the lower bound of transferability is obvious. A model with low accuracy on $\mathcal{D}$ is certainly undesirable for adversarial attacks. Consequently, we minimize this term by training models with sufficient epochs for CIFAR-10 and fine-tuning pre-trained ImageNet classifiers for ImageNette.
@@
It is worth noting that a sensible adversary naturally desires to use more precise surrogates and stronger attacks against unknown targets. However, smoothness and similarity are more complex to understand and optimize compared to fooling probability and natural risk. Therefore, we focus on investigating the connection between smoothness, similarity, and transferability as well as how the various training mechanisms regulate them under the restriction that the other two factors above are fairly acceptable. Prior study [security19] has shown their link with transferability separately, while we explore their joint effect on transferability.
@@ 2.3 Trade-off Under Adversarial Training
Table II: Untargeted and targeted transfer ASRs of AEs crafted against baseline (ST), data augmentations, adversarial training, and gradient regularizations surrogates under the different $L_{\infty}$ budgets (4/255, 8/255, 16/255) on CIFAR-10 using PGD. Results in gray are (more than 0.2%) below the baseline, results in white are close to the baseline ($\pm 0.2\%$). Results in darkest orange are the highest in each respective column, the second darkest are the second highest, and the least dark are others more than 0.2% above the baseline. We conduct experiments using surrogate models trained from 3 different random seeds in each setting and report the average results and the according error bars. In %.
@@
Data distribution shift impairs gradient similarity. Different from model smoothness, it is difficult to understand how exactly AT degrades the gradient similarity. In this work, we attribute this degradation to the data distribution shift induced by the off-manifold examples in AT. Since the emerging of adversarial examples [szegedy2013intriguing], there is a long-held belief that: Clean data lies in a low-dimensional manifold. Even though the adversarial examples are close to the clean data, they lie off the underlying data manifold [ma2018characterizing, song2018pixeldefend, khoury2019on, gilmer2018adversarial, liu2022towards]. Nevertheless, recent researches demonstrate that AEs can also be on-manifold [kwon2021improving, patel2021manifold, Stutz_2019_CVPR, NEURIPS2020_23937b42, liu2022towards], and on-manifold and off-manifold AEs may co-exist [xiao2022understanding]. With this in mind, we analyze how AT enlarges the data distribution shift along with the increment of adversarial budget $\epsilon$.
Given a regular loss function $\ell(f(x),y)$ on the low-dimensional manifold $\mathcal{M}$, the adversarial perturbation $\delta\in\mathbb{R}^{d}$ could make the loss $\ell(f(x+\delta),y)$ sufficiently high with $x+\delta\in\mathcal{M}$ or $x+\delta\notin\mathcal{M}$ [xuICML2022Adversarially2022]. At a high level, AT augments the dataset by adding adversarial examples during each iteration, obtaining an augmented data distribution on $\mathcal{P}_{\mathcal{X}}$, and the adversarial budget $\epsilon$ can be regarded as a parameter that controls the augmentation magnitude. Intuitively, a larger $\epsilon$ induces more space for off-manifold samples, resulting in a larger distribution shift. On the other hand, it is also widely acknowledged that adversarial training with larger perturbation budgets leads to robust overfitting [wu2020adversarial, dong2022exploring, stutz2021relating, pmlr-v119-rice20a]. Accordingly, a larger $\epsilon$ may cause more disbenefit of underfitting to the normal distribution $\mathcal{D}$. Formally, we formalize our hypothesis as follows: